OpenStudy (anonymous):

The decimal number 0.75 means

OpenStudy (solomonzelman):

3 quarters 3/4

OpenStudy (anonymous):

\[0.75=\frac{75}{100}=\frac{25(3)}{25(4)}=\frac{25}{25}\times\frac{3}{4}=\frac{3}{4}\]